OpenStudy (anonymous):

limit of x is 9, find the limit, radical x -3/x-9?? i need step by step please!!!

OpenStudy (anonymous):

\[\lim_{x \rightarrow 9} \sqrt{x}-3/x-9\]

OpenStudy (unklerhaukus):

well simply subsitution wont work, try L'Hôpital's rule

myininaya (myininaya):

\[x-9=(\sqrt{x}+3)(\sqrt{x}-3)\]

myininaya (myininaya):

thats a hint

OpenStudy (unklerhaukus):

or what myininaya said, probably easier for this example

OpenStudy (anonymous):

how did you get to that point tho? radicalx+3 -3 i mean

OpenStudy (dumbcow):

difference of squares a^2 -b^2 = (a+b)(a-b) (sqrtx)^2 - 3^2 = (sqrtx +3)(sqrtx -3)